I am trying to solve a set of equations by the elimination method, but I am having trouble. The set of equations I have to solve are 4x-5y=38 and 7x-8y=22.
Can someone please help?

Answer by nerdybill(7384)

4x-5y=38
7x-8y=22
.
multiply top eq by 7; bottom eq by -4 then ADD:
28x-35y=266
-28x+32y=-88
----------------
-3y = 178
y = 178/(-3)
y = -178/3
y = -59.33
.
find x by substituting above into:
4x-5y=38
4x-5(-59.33)=38
4x=-258.67
x = -64.67